OpenStudy (anonymous):

find a basis for Q[e^(i*pi/3)]

OpenStudy (anonymous):

the basis for this or for the basis for \[\mathbb Z[e^{\frac{\pi}{3}i}]\]? i guess i am confused, because if i recall correctly if you are working over Q it would just be 1 and your root of unity

OpenStudy (anonymous):

sorry, over Q. I got {1, i*sqrt{3}}

OpenStudy (anonymous):

i think that looks good. you definitely need i and root 3, and you get both of those with your basis

OpenStudy (anonymous):

it has been a while, but your extension should be of degree 2, because \[\phi(3)=2\] and you indeed have an extension of degree 2

OpenStudy (anonymous):

Thought that was right. just wasn't sure. Thanks

OpenStudy (anonymous):

sorry i couldn't be more help. are you studying galois?

OpenStudy (anonymous):

Not yet. Not sure we get to that this semester, it may be next semester.

OpenStudy (anonymous):

have fun. it is a great topic

OpenStudy (anonymous):

Thanks, signing off